OpenStudy (anonymous):

A pet show had 32 dogs, 24 cats, and 12 pets of other types. What was the ratio of dogs to all pets at the pet show? A. 8 : 17 B. 8 : 3 C. 8 : 9 D. 4 : 9

OpenStudy (dangerousjesse):

First, tell me the total number of pets :)

OpenStudy (anonymous):

68

OpenStudy (dangerousjesse):

Yep, so what would the initial ratio be of total dogs to total pets?

OpenStudy (anonymous):

what?, im sorry, not to be rude at all, but i didn't get what you are saying

OpenStudy (anonymous):

there are no number in 68 that have to do with the answers

OpenStudy (dangerousjesse):

I know, let's ignore the choices for now. Do you know how to set up a ratio?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

She is trying to say how many dogs are there compared to all of the animals combined.

OpenStudy (anonymous):

32 to 68

OpenStudy (dangerousjesse):

Yes, what Grant said. So, 32 to 68 would be 32:68

OpenStudy (dangerousjesse):

Now we simplify, find the factors of both: \[\large32:1,2,4,8,16\]\[\large62:1,2\]

OpenStudy (anonymous):

In other words, what can both of those numbers simplify down to?

OpenStudy (dangerousjesse):

GCF is 2, I believe, so: \[\large32\div2=16\]\[\large62\div2=31\] 16:31 which is.. none of the choices..

OpenStudy (anonymous):

32 = 8 , 62 = 3?

OpenStudy (anonymous):

so maybe B?

OpenStudy (dangerousjesse):

Where did you get 8 and 3 from?

OpenStudy (anonymous):

im just thinking, cause 62 only simplified to 1, and 2

OpenStudy (anonymous):

32 was 8 and the lowest was 3 on 62's half, so maybe B?

OpenStudy (anonymous):

Remember the total number of pets are 68.

OpenStudy (anonymous):

yeah

OpenStudy (dangerousjesse):

Fml.

OpenStudy (anonymous):

okay wait so my sister is telling me to leave the dogs out and add 24 and 12 instead of getting 68

OpenStudy (anonymous):

so, what do you think?

OpenStudy (dangerousjesse):

I'm sorry, forgot. \[\large32:1,2,4,16\]\[\large68:1,2,4,17\]

OpenStudy (anonymous):

possibly D?

OpenStudy (anonymous):

There we go. LOL

OpenStudy (dangerousjesse):

Divide both by 4 :)

OpenStudy (dangerousjesse):

Shh ;p

OpenStudy (anonymous):

they both made 8

OpenStudy (dangerousjesse):

\[\large32\div4:68\div4\] Try again :)

OpenStudy (anonymous):

Are you sure? I don't think that 68÷4 and 32÷4 are the same answer.

OpenStudy (anonymous):

32 / 4 = 8, 62 / 4 = 15.5

OpenStudy (dangerousjesse):

68, not 62. We're both a little flustered haha.

OpenStudy (anonymous):

Remember that the total number of pets are 68.

OpenStudy (anonymous):

68 / 4 = 17

OpenStudy (anonymous):

A!

OpenStudy (anonymous):

my sister tagged you